40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 3689     Entry time: Mon Oct 11 16:09:10 2010
Author: yuta 
Type: Summary 
Category: SUS 
Subject: current OSEM outputs 

Background:
 The output range of the OSEM is 0-2V.
 So, the OSEM output should fluctuate around 1V.
 If not, we have to modify the position of it.

What I did:
 Measured current outputs of the 5 OSEMs for each 8 suspensions by reading sensor outputs(C1:SUS-XXX_YYPDMON) on medm screens.

Result:

  BS ITMX ITMY PRM SRM MC1 MC2 MC3
UL 1.20 0.62 1.69 1.18 1.74 1.25 0.88 1.07
UR 1.21 0.54 1.50 0.99 1.77 1.64 1.46 0.31
LR 1.39 0.62 0.05 0.64 2.06 1.40 0.31 0.19
LL 1.19 0.88 0.01 0.64 1.64 1.00 0.05 1.03
SD 1.19 0.99 0.97 0.79 1.75 0.71 0.77 0.93

 White: OK (0.8~1.2)
 Yellow: needs to be fixed
 Red: BAD. definitely need fix

ELOG V3.1.3-